Juan invests a total of $28,500 in two accounts. The first account earned a rate of return of 10% (after a year). However, the second account suffered a 7% loss in the same time period. At the end of one year, the total amount of money gained was $-380.00. How much was invested into each account?
$ was invested at 10% and
$ was invested at 7%.

Found 2 solutions by JoelSchwartz, Boreal:
Answer by JoelSchwartz(130)

x=amount invested with a return rate of 10 percent
y=amount invested with a loss of seven percent
x+y=28,500
x=28500-y
.1x-.07y=-380
.1(28500-y)-.07y=-380
2850-.1y-.07y=-380
-.17y=-3230
y=19,000
x=9,500

Answer by Boreal(15235)

amount into first =x
amount into second=28,500-x
.10x - .07(28500-x)=-380
.10x-1995+.07x= -380
.17x= 1615
x=9500; 10% interest gives + $950.
28500-x=19000
19000*-0.07= -1330
Difference is -$380